Now we are going to look at the problem of an object sliding down a frictionless inclined plane. Suppose a toboggan loaded with vacationing students (total weight w) slides down a long, snow-covered slope. The hill slopes at a constant angle α, and the toboggan is so well waxed that there is virtually no friction. Find the toboggan’s acceleration and the magnitude n of the normal force the hill exerts on the toboggan.

At what angle does the hill slope if the acceleration is g/4?
Answer #1

Now ,

on a hill with angle alpha

acceelration is given as

a = g * sin(alpha)

as a = g/4

g/4 = g * sin(alpha)

1/4 = sin(alpha)

alpha = 14.5 degree

the angle of hill slope is 14.5 degree